This year Riverside began work on an outdoor amphitheater and concession stand at the city’s park. It is to be financed by a $5,300,000 bond issue and supplemented by a $680,000 General Fund transfer. The following transactions occurred during the current year:

  1. The General Fund transferred $680,000 to the Park Building Capital Projects Fund.
  2. A contract was signed with Restin Construction Company for the major part of the project on a bid of $4,500,000.
  3. Preliminary planning and engineering costs of $87,000 were vouchered for the Great Pacific Engineering Company. (This cost had not been encumbered.)
  4. A payable was recorded for an $27,500 billing from the Water and Sewer enterprise fund for the cost of extending water pipes to the new concession stand.
  5. An invoice in the amount of $1,900,000 was received from Restin for progress to date on the project.
  6. The $5,300,000 bonds were issued at par.
  7. The amount billed by the contractor (see Transaction 5) less 5 percent retained was paid.
  8. Temporary investments were purchased at a cost of $1,980,000.
  9. Closing entries were prepared as of year-end. Assume that $3,400,000 of the encumbrances outstanding will be paid from the bond proceeds and are therefore considered restricted resources. The remaining fund balance is assigned.

Prepare a balance sheet for the Park Building Capital Projects Fund as of year-end.

Prepare a statement of revenues, expenditures, and changes in fund balances for the period.
Answer #1

RIVERSIDE

Park Building Capital Projects Fund

Balance Sheet, December 31

Assets

Cash

2195000

Investments

1980000

Total assets

4175000

Liabilities and fund balances

Liabilities:

Contracts Payable—Retained Percentage

95000

Due to Other Funds

27500

Vouchers Payable

87000

Total Liabilities

209500

Fund Balances:

Fund Balance—Restricted

3400000

Fund Balance—Assigned

565500

Total Fund Balances

3965500

Total Liabilities and Fund Balances

4175000

Cash = 680000+5300000-(1900000*0.95)-1980000 = 7785000

19000000*0.05 = 95000

5300000-3400000-(27500+87000+1900000)

RIVERSIDE

Park Building Capital Projects Fund

Statement Of Revenues, Expenditures, and Changes in Fund Balances

For the Period Ended December 31

Expenditures:

Construction Expenditures

2014500

Other Financing Sources and Uses:

Interfund Transfers In

680000

Proceeds of Bonds Sold

5300000

5980000

Excess of Revenues and Other Sources Over Expenditures

3965500

Fund Balances, January 1

0

Fund Balances, December 31

3965500